OpenStudy (anonymous):

A store charges their customers based on the number of items that they buy. Jane bought four items and was charged $15 and George bought twelve items and was charged $39. Write a function to model the cost based on the number of items you buy. Then determine how much you are charged if you purchase 22 items.

OpenStudy (hitaro9):

Okay. So why don't we set up some equations, since the answer doesn't seem readily apparent.

OpenStudy (hitaro9):

Let's let x = the cost per item

OpenStudy (hitaro9):

So 15 = x(12) since jane bought 4 items.

OpenStudy (hitaro9):

Oh wait, 15 = x(4) my bad

OpenStudy (hitaro9):

and 39 = x(12)

OpenStudy (hitaro9):

So, if we divide these out we find that the two do not yield the same result.

OpenStudy (hitaro9):

So there must be some added cost, y.

OpenStudy (hitaro9):

So we set up the equations to be 15 = 4x + y 39 = 12x+y

OpenStudy (hitaro9):

There you can either do substitution, or subtraction. I'm not sure which your teacher taught you?

OpenStudy (anonymous):

the answer is 10